OpenStudy (anonymous):

Desperately need help with this Imagine you are the owner of a pizza shop that offers three different types of crusts, seven different toppings, and three different cheeses.Create an advertisement for your shop that lists all the choices you offer and advertises a three-topping pizza with one type of cheese for $9.99. Complete this statement and include it on your advertisement: "We offer _______ different 3-topping pizza choices for you to choose from!" How do I find the number of choices?

OpenStudy (anonymous):

Choose a crust: \(^3C_1\) Choose three toppings: \(^7C_3\) Choose a cheese \(^3C_1\) Multiple them all together.

OpenStudy (anonymous):

what does C mean?

OpenStudy (kc_kennylau):

\[\Large C^n_r=\frac{n!}{(n-r)!r!}\]

OpenStudy (kc_kennylau):

I don't know why he put the first number in front of the C but they mean the same thing don't they

OpenStudy (kc_kennylau):

C stands for combination

OpenStudy (anonymous):

So it's the formula?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

what does r and n mean?

OpenStudy (kc_kennylau):

Just two variables in a formula

OpenStudy (kc_kennylau):

Wait I think C stands for choose

OpenStudy (kc_kennylau):

Because we pronounce \(\Large C^n_r\) as "n choose r".

OpenStudy (kc_kennylau):

That means the number of ways to choose "r" things from "n" things

OpenStudy (kc_kennylau):

ignoring order

OpenStudy (anonymous):

So you create a formula for each right?

OpenStudy (kc_kennylau):

There's only one formula \(\Large C^n_r=\frac{n!}{(n-r)!r!}\)